OpenStudy (anonymous):

Solve the quadratic equation y2 - 7y - 18 = 0 by factoring.

OpenStudy (anonymous):

then (x - 2) = 0 and (x + 9) = 0 x = 2 x = -9

OpenStudy (anonymous):

what 2 factors add to -7 and multiply to - 18? -9 and 2 ? yes ? (x + 2)(x - 9)

OpenStudy (anonymous):

\[y ^{2} - 7y -18 = 0\]\[(y+2)(y-9)=0\]\[y=2 , y=-9\]